### v1

@@ 3.4 Adapting ViT to Task-specific FSAR
FSAR studies how to classify unseen class query videos with only a limited number of labeled samples available. Following the episodic training, we randomly sample a $C$-way $K$-shot task, which has $C$ action classes and $K$ labeled samples in each class, from the training set $\mathcal{D}_{tr}$ for each episode during training. All the ${C\times K}$ video samples form the support set $\mathcal{S}=\{x_{i}\}_{i=1}^{CK}$. Additionally, there is a query set $\mathcal{Q}=\{x_{i}\}_{i=1}^{Q}$ from the same $C$ action classes. The learning objective of each task is to classify each query video $q\in\mathcal{Q}$ into one of the $C$ classes. Although the temporal adaptation proposed by AIM gives vanilla ViT the ability of temporal modeling, it remains unaddressed that the ViT model could not attend to different videos and enhance the task-specific discriminative features.
Next, we describe how we empower the frozen ViT with the ability to adapt embeddings with task-specific information for few-shot action recognition. As described in Section 3.3, we first map all the videos in the whole sampled task (both support set and query set) into their patch embeddings by Eq. (1) and Eq. (2), then we get the input features of the first ViT block $\mathcal{S}_{0}=[x_{0}^{1},...,x_{0}^{CK}]$ and $\mathcal{Q}_{0}=[x_{0}^{1},...,x_{0}^{Q}]$, where $x_{0}^{i}\in\mathbb{R}^{T\times(N+1)\times D}$ denotes the patch embeddings of the $i$-th video both for the support set and query set. Then, these patch embeddings are fed into several frozen ViT blocks to exchange spatial information between different tokens. It should be noted that we only introduce tunable adapters in the last L ViT blocks, which is a choice based on empirical study. We will further analyze the impact of this decision in the ablation experiment. Starting from the (N-L+1)-th layer, we introduce tunable adapter layers into the original model. Specifically, besides T-MSA and S-MSA, we further reuse the frozen MSA layer as Task-MSA after the temporal and spatial adaptation to perform task-specific self-attention across the tokens at the same spatial-temporal location in different videos (see Fig. 4). Take the support features as an example:
@@
$\mathcal{S}_{l}$ $={\rm Adapter}({\rm LN}(\mathcal{S}_{l}^{task}))$
$+{\rm MLP}({\rm LN}(\mathcal{S}_{l}^{task}))+\mathcal{S}_{l}^{task}$ (9)
@@
Finally, we concatenate the [class] tokens of each frame from the last ViT block to get the support set video representations $\mathcal{F}_{\mathcal{S}}\in\mathbb{R}^{CK\times T\times D}$ and the query set video representations $\mathcal{F}_{\mathcal{Q}}\in\mathbb{R}^{Q\times T\times D}$. Both the above prototypes will be used to calculate matching score in the metric step.
@@ 3.5 Semantic Order Adaptation
Essentially, our primary task is to construct a semantic corpus that contains dynamic descriptions of different actions. We call the LLM to generate fine-grained sub-actions by the temporal sequence with the following prompts:" Given an action label {LABEL}, you should describe three states of each action." For example, if the input label is long jump, the corresponding output is a list followed ["Run up for momentum.", "Take off with a powerful leap.", "Land and maintain balance."]
After acquiring the fine-grained descriptions above, we concatenate the original label with the three sub-actions to form three prompts, which are then used as input for the semantic branch of CLIP. Such as "A video of action about {label}: {sub-actions}".
Inspired by PEFT, we froze the text encoder and a tunable adapter is added after the last $M$-layer to achieve semantic order adaptation. We reuse the frozen MSA layers as O-MSA, which could model the 3 sub-actions. For instance:
@@
where the output semantic feature of class $c$ in $j$-th ViT blockthe ($j\in\{N-M+1,N-M+2,...,N\}$) can be described as $\mathcal{T}_{j}^{c}=\{\mathcal{T}_{j}^{c_{1}},\mathcal{T}_{j}^{c_{2}},\mathcal{T}_{j}^{c_{3}}\}\in\mathbb{R}^{77\times 3\times D}$, where $77$ presents the context length in text encoder. After the whole branch, we concatenate the features of different actions within a task to obtain a vector $\mathcal{T}_{j}^{C}\in\mathbb{R}^{C\times 77\times 3\times D}$
Finally, we select the [class] token from semantic features $\mathcal{F}_{\mathcal{T}}\in\mathbb{R}^{C\times 3\times D}$ as the output followed by CLIP from the last layer of ViT block.
Figure 6: The structure of our alignment module. We use cross-attention as the alignment module. The ($i\mbox{-}1$)-th frame serves as $q$ and the $i$-th frame serves as $k$ and $v$.
3.6 Cross modal Alignment
Generally, considering that video has richer information and higher matching score than text, aligning video features to text is good choice. As shown in Fig. 6, we propose a novel alignment method that model the information between adjacent frames. The algorithm could be described as:
@@ 3.6 Cross modal Alignment
where $\mathcal{F}_{Q}^{i}\in\mathbb{R}^{C\times D}$ means the $i$-th frame of the input visual information, and for ease of understanding we define the aligned features as ¯$\bar{\mathcal{F}}_{Q}$.
The final question is how to calculate the classification result. The result is composed by two parts: video matching score of video prototypes and video-text matching score. For the support video feature $\mathcal{F}_{\mathcal{S}}$ and query video feature $\mathcal{F}_{\mathcal{Q}}$, the support video prototype can be denoted as:
where ^$\hat{\mathcal{F}}_{\mathcal{S}}^{c}\in\mathbb{R}^{C\times T\times D}$ denotes the class prototype of the $c$-th support class. We directly average all the support video features as the class prototype for each class. To calculate the classification probability, we measure the distance between the query features and the class prototypes by an arbitrary metric module, e.g., TRX [30], OTAM [4] and Bi-MHM [43], which can formulated by:
where $\mathcal{P}^{c}_{i}$ is the probability of the $c$-th class for the $i$-th query video in query set and ^$\hat{\mathcal{F}}_{\mathcal{S}}=[\hat{\mathcal{F}}_{\mathcal{S}}^{1},\hat{\mathcal{F}}_{\mathcal{S}}^{2},...,\hat{\mathcal{F}}_{\mathcal{S}}^{c}]\in\mathbb{R}^{C\times T\times D}$ stands for all the class prototypes of the whole support set. For the metric modules, we will demonstrate in our experiments that our proposed Task-Adapter++ consistently improves the performance of any metric methods. Besides that, owning to the superiority of our task-specific adaptation during feature extraction, excellent performance can be achieved simply by averaging the frame-to-frame cosine similarity without complex metric measurements.
We calculate the possibility distribution of cross-modality by cosine similarity using semantic features $\mathcal{F}_{\mathcal{T}}$ and aligned query features ¯$\bar{\mathcal{F}}_{\mathcal{Q}}$.
$\dot{\mathcal{P}}=\sum_{s=1}^{3}\left[\begin{array}[]{c}\mathcal{F}_{\mathcal{T}}^{1s}\ \mathcal{F}_{\mathcal{T}}^{2s}\ \cdots\ \mathcal{F}_{\mathcal{T}}^{Cs}\end{array}\right]\left[\begin{array}[]{c}\bar{\mathcal{F}}_{\mathcal{Q}_{1}}^{[2s-1]}+\bar{\mathcal{F}}_{\mathcal{Q}_{1}}^{[2s]}+\bar{\mathcal{F}}_{\mathcal{Q}_{1}}^{[2s+1]}\\ \bar{\mathcal{F}}_{\mathcal{Q}_{2}}^{[2s-1]}+\bar{\mathcal{F}}_{\mathcal{Q}_{2}}^{[2s]}+\bar{\mathcal{F}}_{\mathcal{Q}_{2}}^{[2s+1]}\\ \vdots\\ \bar{\mathcal{F}}_{\mathcal{Q}_{C}}^{[2s-1]}+\bar{\mathcal{F}}_{\mathcal{Q}_{C}}^{[2s]}+\bar{\mathcal{F}}_{\mathcal{Q}_{C}}^{[2s+1]}\\ \end{array}\right]$ (16)
@@
where $\mathcal{F}_{\mathcal{T}}^{Cs}$ means the $s$-th sub-action features of class $C$, and ¯$\bar{\mathcal{F}}_{\mathcal{Q}_{C}}^{[2s]}$ means the $2s$-th frame aligned query feature in class $C$. The process is explained well by Fig. 7.
@@
Figure 7: The illustration of matching process by stages. The aligned query features are divided into three segments and matched with the features of three sub-actions respectively.
Different from previous works, we cut query features into three pieces instead of simply averaging them. Therefore, the matching between query videos and sub-actions can be coupled in temporal dimension. We combine the zero-shot results and few-shot results in the same way as CLIP-FSAR [41] which uses a simple element-wise multiplication. Thus, the final score is calculated in the form of a product:
$\text{Score}=\mathcal{P}\times\dot{\mathcal{P}}/9$ (17)
@@
During the training stage, classification probabilities for each task serve as the logits for computing cross-entropy loss and backpropagating gradients. Importantly, we solely fine-tune the parameters of the additional adapter layers while keeping the original pre-trained model frozen. In testing, we freeze the entire backbone, including the adapters, and directly extract features for the unseen test classes.
@@
Table 1: This part reported the results focused on temporal-related dataset (SSv2-Small and SSv2-Full). Comparison with current SOTA few-shot action recognition methods on 5-way 1-shot, 5-way 3-shot and 5-way 5-shot benchmarks. The best results are highlighted in bold.
@@
[…]
@@
MA-CLIP [51] ArXiv’23 ViT-B CLIP PEFT 59.1 - 64.5 63.3 - 72.3
TSAM [17] Arxiv’24 ViT-B CLIP PRFT 60.5 - 66.7 65.8 - 74.6
@@ 4.1 Experimental Settings
Datasets.We evaluate our method on four datasets that widely used in few-shot action recognition, namely HMDB51 [15], UCF101 [34], Kinetics [5] and SSv2 [11]. The first three datasets focus on scene understanding, while the last dataset, SSv2, has been shown to require more challenging temporal modeling [38, 4]. Particularly, We use the few-shot split proposed by [58] for HMDB51 and UCF101, with 31/10/10 classes and 70/10/21 classes for train/val/test, respectively. Following [64], Kinetics is used by selecting a subset which consists of 64, 12 and 24 training, validation and testing classes. For SSv2, we provide the results of two publicly available few-shot split, i.e., SSv2-Small [65] and SSv2-Full [4]. Both SSv2-Small and SSv2-Full contain 100 classes selected from the original dataset, with 64/12/24 classes for train/val/test, while SSv2-Full contains 10x more videos per class in the training set.
Implementation details. As previously illustrated, on one hand we aim to adapt the image pre-trained ViT model for task-specific few-shot action recognition, on the other hand we use order adaptation to model semantic information of sub-actions. To better demonstrate the model’s capabilities, we utilized a CLIP pre-trained ViT model [31] as the backbone. For a fair comparison with the existing works [41, 51], we uniformly sample 8 frames for each video, and crop the frames to 224$\times$224 as the input resolution. Before the training, we use GPT-4 [28] as LLM to enhance our label information. The semantic corpus is saved as json file without any manual modification. During the training stage, random crop is used for data augmentation. We freeze the original pre-trained model and only fine-tune the introduced adapters (as shown in Fig. 2 (c)) using the SGD optimizer with a learning rate of 0.001. For the testing stage, we freeze all the parameters of our backbone to extract task-specific discriminative video features and semantic features for measuring the distances. All reported results in our paper are the average accuracy over 10,000 few-shot action recognition tasks.
Table 2: This part reported the results focused on scene-related datasets (including HMDB51, UCF101, and Kinetics). Comparison with current SOTA few-shot action recognition methods on 5-way 1-shot, 5-way 3-shot and 5-way 5-shot benchmarks. The best results are highlighted in bold.
@@
[…]
@@
Task-Adapter++ - ViT-B CLIP PEFT 86.5 90.6 92.0 98.9 99.5 99.7 96.0 97.3 97.5
@@ 4.2 Comparison with state-of-the-art methods
Table 1 and Table 2 presents a comprehensive comparison of our method with recent few-shot action recognition methods. As our Task-Adapter++ is designed for ViT-based architectures, we use CLIP [31] pre-trained ViT to make a fair comparison with recent few-shot action recognition works that use large-scale pretrained models. Of particular note, by comparing the results of works that use CLIP pre-trained models and those that use ImageNet (IN-21K) pre-trained models, we can conclude that the performance of few-shot action recognition indeed benefits from large-scale pretraining, especially for scene-related datasets.
As shown in Table 1, our Task-Adapter++ can achieve best results on SSv2-Small and SSv2-Full, indicating that our method can model the temporal information well. For Average 1-shot, 3-shot and 5-shot, our method outperforms the exsisting SOTA methods by 3.1% on SSv2-Small and 1.4% on SSv2-Full, respectively. Compared with full fine-tuning method CLIP-FSAR [41], we can see our method has an overall improvement (from 3.4% to 11.4%). Compared with PEFT method TSAM [17], our method also has an overrall improvement (from 0.9% to 7.0%) on two datasets. We generally believe that SSv2 requires more temporal modeling than other three datasets, making it much more challenging, and we have achieved a significant improvement of at least 3% on SSv2-Small compared to the SOTA, indicating that our task can better adapt to the classification of complex actions.
In addition, we display the results on HMDB51, UCF101 and Kinetics in Table 2. It is obvious that we achieve SOTA results than other methods. Compared with full fine-tuning method CLIP-FSAR [41], we achieve an average improvement of 6.9% on HMDB51, 1.3% on UCF101 and 1.7% on Kinetics. Compared with PEFT method TSAM [17], our method also has an average improvement of 2.6% on HMDB51, 0.5% on UCF101 and 0.1% on Kinetics. These scene-related datasets require strong background understand ability which means our model can handle complex scene knowledge well.
Table 3: We choose frozen backbone as the baseline and compare the performance of different settings for 5-way 1-shot task on SSv2-Small.
@@
Table 4: We choose our Task-Adapter as the baseline and compare the performance of different settings for 5-way 1-shot task on SSv2-Small.
Task-Adapter Three Sub-actions Order Adaptation Cross Attention Alignment Acc

### v2

@@ 3.4 Adapting ViT to Task-specific FSAR
FSAR studies how to classify unseen class query videos with only a limited number of labeled samples available. Although the temporal adaptation proposed by AIM gives vanilla ViT the ability of temporal modeling, it remains unaddressed that the ViT model could not attend to different videos and enhance the task-specific discriminative features.
Therefore, we describe how we empower the frozen ViT with the ability to adapt embeddings with task-specific information for few-shot action recognition. As described in Section 3.3, we first map all the videos in the whole sampled task (both support set and query set) into their patch embeddings by Eq. (1) and Eq. (2), then we get the input features of the first ViT block $\mathcal{S}_{0}=[x_{0}^{1},...,x_{0}^{CK}]$ and $\mathcal{Q}_{0}=[x_{0}^{1},...,x_{0}^{Q}]$, where $x_{0}^{i}\in\mathbb{R}^{T\times(N+1)\times D}$ denotes the patch embeddings of the $i$-th video both for the support set and query set. Then, these patch embeddings are fed into several frozen ViT blocks to exchange spatial information between different tokens. It should be noted that we only introduce tunable adapters in the last L ViT blocks, which is a choice based on empirical study. We will further analyze the impact of this decision in the ablation experiment. Starting from the (N-L+1)-th layer, we introduce tunable adapter layers into the original model. Specifically, besides T-MSA and S-MSA, we further reuse the frozen MSA layer as Task-MSA after the temporal and spatial adaptation to perform task-specific self-attention across the tokens at the same spatio-temporal location in different videos (see Fig. 3). Take the support features as an example:
@@
$\mathcal{S}_{l}$ $={\rm Adapter}({\rm LN}(\mathcal{S}_{l}^{task}))+{\rm MLP}({\rm LN}(\mathcal{S}_{l}^{task}))+\mathcal{S}_{l}^{task}$ (9)
@@
Finally, we concatenate the [class] tokens of each frame from the last ViT block to get the support set video representations $\mathcal{F}_{\mathcal{S}}\in\mathbb{R}^{CK\times T\times D}$ and the query set video representations $\mathcal{F}_{\mathcal{Q}}\in\mathbb{R}^{Q\times T\times D}$. The video metric score is calculated based on the support prototypes and query features.
For the support video features $\mathcal{F}_{\mathcal{S}}$, the support video prototype can be denoted as:
where ^$\hat{\mathcal{F}}_{\mathcal{S}}^{c}\in\mathbb{R}^{T\times D}$ denotes the class prototype of the $c$-th support class. We directly average all the support video features as the class prototype for each class.
To calculate the classification probability, we measure the distance between the query features and the class prototypes by an arbitrary metric module, e.g., TRX [15], OTAM [14] and Bi-MHM [16], which can formulated by:
where $\mathcal{P}^{c}_{i}$ is the probability of the $c$-th class for the $i$-th query video in query set and ^$\hat{\mathcal{F}}_{\mathcal{S}}=[\hat{\mathcal{F}}_{\mathcal{S}}^{1},\hat{\mathcal{F}}_{\mathcal{S}}^{2},...,\hat{\mathcal{F}}_{\mathcal{S}}^{c}]\in\mathbb{R}^{C\times T\times D}$ stands for all the class prototypes of the whole support set.
In our experiments of different metric modules, we will demonstrate that our proposed Task-Adapter++ consistently improves the performance of any metric methods. Besides that, owning to the superiority of our task-specific adaptation during feature extraction, excellent performance can be achieved simply by averaging the frame-to-frame cosine similarity without complex metric measurements.
Figure 5: Illustration of order adaptation in semantic branch. For semantic order adaptation, we introduce semantic order adapter after MSA to model the sequential information of sub-actions. This part illustrates the computational process of the text encoder given a 2-way 1-shot learning task.
@@ 3.5 Semantic Order Adaptation
Considering that existing works overlook the fine-grained sequential modeling in the semantic feature extraction process, we propose the semantic order adaptation. Essentially, our primary task is to construct a semantic corpus that contains dynamic descriptions of different actions. We prompt the LLMs to generate fine-grained sub-actions with following instruction: “Given an action label {LABEL}, describe three stages of the action.” The LLMs generate corresponding descriptions of the beginning, process, and end for each input action label. For example, if the input label is “Long Jump”, the corresponding output is a list delineates the beginning, process, and end of the action as follows: [“Run up for momentum.”, “Take off with a powerful leap.”, “Land and maintain balance.”]
After obtaining the fine-grained descriptions as above, we concatenate the original label with the three sub-actions respectively to form three prompts, which are then used as input for the semantic branch of CLIP, such as “A video of action about {LABEL}: {SUB-ACTIONS-1}”, “A video of action about {LABEL}: {SUB-ACTIONS-2}” and “A video of action about {LABEL}: {SUB-ACTIONS-3}”.
As shown in Fig. 5, to enhance the representation of sequential relationships between sub-actions, inspired by PEFT, we freeze the text encoder and introduce a tunable adapter after the last $M$-layer to achieve semantic order adaptation. Specifically, we reuse the frozen MSA layers as O-MSA, which could model the sequential information across the three sub-actions as follows:
@@
where the output semantic feature of class $c$ in the $j$-th ViT block ($j\in\{N-M+1,N-M+2,...,N\}$) is denoted as $\mathcal{T}_{j}^{c}=\{\mathcal{T}_{j}^{c_{1}},\mathcal{T}_{j}^{c_{2}},\mathcal{T}_{j}^{c_{3}}\}\in\mathbb{R}^{l\times 3\times D}$, and $l$ presents the sequence length in text encoder.
Finally, for each action within a task, we select the end-of-text token from the text features in the last block and concatenate them to form the output semantic feature for cross-modal matching.
Figure 6: The structure of our fine-grained cross-modal alignment. We use cross-attention for fine-grained temporal modeling, where the ($i\mbox{-}1$)-th frame serves as $q$ and the $i$-th frame serves as $k$ and $v$.
3.6 Fine-grained Cross-modal Alignment
Considering that videos contain richer spatio-temporal information while semantics are more abstract, mapping videos into the semantic space aligns more naturally with the principles of information extraction. Thus, we propose a fine-grained cross-modal alignment framework that models the temporal relationships between adjacent frames. In this framework, visual features are aligned to correspond with the same temporal stages as their semantic descriptions. As illustrated in Fig. 6, our method enables more precise alignment by modeling frame-to-frame dynamics in the temporal context with cross-attention:
@@ 3.6 Fine-grained Cross-modal Alignment
where $\mathcal{F}_{Q}^{i}\in\mathbb{R}^{Q\times D}$ denotes the $i$-th frame of the query’s visual feature set, and for ease of understanding we define the aligned features as ¯$\bar{\mathcal{F}}_{Q}\in\mathbb{R}^{Q\times(T-1)\times D}$. By default, $T$ is set to 8. As no padding is applied along the temporal dimension, the length of the aligned features is reduced by one compared to the original input sequence..
For the $q$-th aligned visual features ¯$\bar{\mathcal{F}}_{\mathcal{Q}_{q}}$ in the query set, we divide it into three temporal segments, corresponding to the three different sub-actions, referred to as:
¯$\bar{\mathcal{F}}_{\mathcal{Q}_{q}}=[\frac{1}{3}\sum^{3}_{i=1}\bar{\mathcal{F}}_{\mathcal{Q}_{q}}^{i},\frac{1}{3}\sum^{5}_{i=3}\bar{\mathcal{F}}_{\mathcal{Q}_{q}}^{i},\frac{1}{3}\sum^{7}_{i=5}\bar{\mathcal{F}}_{\mathcal{Q}_{q}}^{i}]\in\mathbb{R}^{3\times D}$ (16)
Then, the probability distribution is calculated based on the cosine similarity between the semantic feature $\mathcal{F}_{\mathcal{T}}$ and the aligned query feature segment ¯$\bar{\mathcal{F}}_{\mathcal{Q}_{q}}$, which can be formulated as:
$\dot{\mathcal{P}}_{q}^{c}=\frac{1}{3}\sum_{s=1}^{3}\text{Cos}(\bar{\mathcal{F}}_{\mathcal{Q}_{q}}[s],\mathcal{F}_{\mathcal{T}}^{c}[s])$ (17)
@@
where $\dot{\mathcal{P}}_{q}^{c}$ is the probability of the $c$-th class for the $q$-th video in query set. ¯$\bar{\mathcal{F}}_{\mathcal{Q}_{q}}[s]$ means the $s$-th query feature segment for the $q$-th query video and $\mathcal{F}_{\mathcal{T}}^{c}[s]$ means the $s$-th sub-action features of class $c$. The process is also illustrated in Fig. 7.
@@
Figure 7: The stage-wise illustration of the cross-modal feature matching process. The aligned query features are divided into three segments and matched with the features of three sub-actions respectively.
Notably, instead of averaging temporal video features for global matching, we divide them into three temporal segments to facilitate fine-grained cross-modal alignment. This design enables the visual and textual features to be matched at each corresponding stage individually, resulting in more precise and discriminative matching result.
3.7 Calculation of Classification Results
The final classification score consists of two components: the video matching score $\mathcal{P}$, as demonstrated in Section 3.4, which measures the similarity between support and query features based on video prototypes; and the video-text matching score $\mathcal{\dot{P}}$, introduced in Section 3.6, which evaluates the alignment between semantic features and query features. These two scores are combined via element-wise multiplication to jointly consider both visual and semantic similarities. Thus, the final score is calculated in the form of a product:
$\text{Score}=\mathcal{P}\times\dot{\mathcal{P}}$ (18)
@@ 3.7 Calculation of Classification Results
The similarity scores for each task are treated as logits and used to compute the cross-entropy loss, enabling gradient backpropagation during training. Notably, we only fine-tune the parameters of the additional adapter layers, while keeping the backbone of the pre-trained model frozen. During testing, the entire model, including the adapters, is fixed, and the highest-scoring class for each query video is selected as the final prediction.
@@
Table 1: Comparison with current SOTA few-shot action recognition methods on temporal-related datasets (SSv2-Small and SSv2-Full) with 5-way 1-shot, 5-way 3-shot and 5-way 5-shot settings. The best results are highlighted in bold.
@@
MASTAF [59] WACV’23 ViViT JFT Full Fine-tuning 45.6 - - 60.7 - -
@@
MA-FSAR [19] ArXiv’23 ViT-B CLIP PEFT 59.1 - 64.5 63.3 - 72.3
TSAM [21] AAAI’25 ViT-B CLIP PEFT 60.5 - 66.7 65.8 - 74.6
@@ 4.1 Experimental Settings
Datasets. We evaluate our method on five datasets that widely used in few-shot action recognition, namely HMDB51  [61], UCF101 [62], Kinetics [7] and SSv2 [8]. The first three datasets focus on scene understanding, while the last dataset, SSv2, has been shown to require more challenging temporal modeling [63, 14]. Particularly, We use the few-shot split proposed by [43] for HMDB51 and UCF101, with 31/10/10 classes and 70/10/21 classes for train/val/test, respectively. Following [54], Kinetics is used by selecting a subset which consists of 64, 12 and 24 training, validation and testing classes. For SSv2, we provide the results of two publicly available few-shot split, i.e., SSv2-Small [64] and SSv2-Full [14]. Both SSv2-Small and SSv2-Full contain 100 classes selected from the original dataset, with 64/12/24 classes for train/val/test, while SSv2-Full contains 10x more videos per class in the training set.
Implementation details. As previously illustrated, on one hand we aim to adapt the image pre-trained ViT model for task-specific few-shot action recognition, on the other hand we use order adaptation to model semantic information of sub-actions. To better demonstrate the model’s capabilities, we utilize a CLIP pre-trained ViT model [1] as the backbone. For a fair comparison with the existing works [60, 19], we uniformly sample 8 frames for each video, and crop the frames to 224$\times$224 as the input resolution. Before the training, we use GPT-4 [65] to enhance our label information. The semantic corpus is saved as json file without any manual modification. During the training stage, random crop is used for data augmentation. We freeze the original pre-trained model and only fine-tune the introduced Task adapters and order adapters using the SGD optimizer with a learning rate of 0.001. All reported results in our paper are the average accuracy over 10,000 few-shot action recognition tasks.
Table 2: Comparison with current SOTA few-shot action recognition methods on scene-related datasets (HMDB51, UCF101 and Kinetics) with 5-way 1-shot, 5-way 3-shot and 5-way 5-shot settings. The best results are highlighted in bold.
@@
MASTAF [59] WACV’23 ViViT JFT Full Fine-tuning 69.5 - - 91.6 - - - - -
@@
Task-Adapter++ - ViT-B CLIP PEFT 86.5 90.6 92.0 98.9 99.5 99.7 96.2 97.3 97.5
@@ 4.2 Comparison with state-of-the-art methods
Table 1 and Table 2 presents a comprehensive comparison of our method with recent few-shot action recognition methods on both temporal-related datasets and scene-related datasets. As our Task-Adapter++ is designed for ViT-based architectures, we use CLIP [1] pre-trained ViT to make a fair comparison with recent few-shot action recognition works that use large-scale pretrained models. Of particular note, by comparing the results of works that use CLIP pre-trained models and those that use ImageNet (IN-21K) pre-trained models, we can conclude that the performance of few-shot action recognition indeed benefits from large-scale pretraining, especially for scene-related datasets.
As shown in Table 1, our Task-Adapter++ can achieve the best results on SSv2-Small and SSv2-Full, leading previous SOTA method with a large margin, which indicates the superiority of our method on temporal challenging few-shot learning tasks. The results of Task-Adapter compared with MA-FSAR [19] on the SSv2 datasets demonstrate the effectiveness of our proposed method for task-specific adaptation, showing a significant improvement of 5.7% for the 5-shot task on SSv2-Small and 8.0% for the 1-shot task on SSv2-Full, respectively. Our Task-Adapter have achieved an average improvement of 3.4% on SSv2-Small and 4.9% on SSv2-Full compared to the SOTA method MA-FSAR [19]. Our Task-Adapter++ have achieved further improvement on the basis of Task-Adapter, specifically improving 3.1% on SSv2-Small and 1.4% on SSv2-Full, benefiting from the semantic order adaptation and fine-grained cross-modal matching ability. Compared with full fine-tuning method CLIP-FSAR [60], we can see our Task-Adapter++ has a significant overall improvement (from 3.4% to 11.4%) with fewer learnable parameters (7.5 M vs. 89.3 M). Compared with PEFT method TSAM [21], our Task-Adapter++ also has a comprehensive improvement (from 0.9% to 7.0%) on SSv2. Our Task-Adapter++ has further expanded the average improvement, 4.3% on SSv2-Small and 3.9% on SSv2-Full compared to the SOTA method TSAM [21], indicating that our method can better adapt to the temporal complex FSAR tasks with dual adaptation.
In addition, we report the results on HMDB51, UCF101 and Kinetics in Table 2. Especially for 5-shot tasks, our Task-Adapter outperforms the existing SOTA methods MA-FSAR [19] by 0.9% on HMDB51, 0.2% on UCF101 and 0.8% on Kinetics, respectively. Above our Task-Adapter, Task-Adapter++ achieve higher score by 2.9% on HMDB51, 0.7% on UCF101 and 1.0% on Kinetics, respectively. Compared with full fine-tuning method CLIP-FSAR [60], our Task-Adapter++ achieve an average improvement of 6.9% on HMDB51, 1.3% on UCF101 and 1.8% on Kinetics. Compared with PEFT method TSAM [21], our method also has an average improvement of 2.6% on HMDB51, 0.5% on UCF101 and 0.2% on Kinetics. These scene-related datasets require strong background understand ability which means our model can handle complex scene knowledge well.
Table 3: Ablation study of Task-adapter. We choose frozen backbone as the baseline under 5-way 1-shot settings on SSv2-Small.
@@
Table 4: Ablation study of Task-adapter++. Task-Adapter is chosen as the baseline under 5-way 1-shot settings on SSv2-Small.
Task-Adapter Three Sub-actions Order Adaptation Cross-modal Alignment Acc